LBC on 14 September 2009 UT

Clouds.

05:10 We open dome (still some thunderstorms south of Wilcox) T_air=9.4 RH=80% T_mirr=11.8

Right Shutter alarm! Left Opened.

05:40 Shutter problem fixed

05:42 Slew to WT10_238 to check Pointing (AZ=179 EL=+66)

Old CA=-28.96 IE=-23.17 New CA=-100.2 IE=+67.8

05:54 Problem with IDL licence on lbtmu42; dofpia doesn't work.Norm on call 06:06 Problem with IDL fixed

06:15 T_air = 7.9 now (dropped by 4^C since begin of night)

DOFPIA took 7 iterations to converge. RED struggled with Z8 Mirror now 4^C warmer than air !

LBCB

(Z4) (Z5) (Z6) (Z7) (Z8) (Z11) (Z22)| seeing WIQ


-5 352 -69 -16 78 -69 -19 1.16 0.14

LBCR

(Z4) (Z5) (Z6) (Z7) (Z8) (Z11) (Z22)| seeing WIQ


-35 -351 208 -213 -152 -60 -1 0.97 0.19

06:19 Fast OB VV to test Tech chips

No Zernike polymonials appeared on the LBC log ! They should have.

06:30 Vincenzo put an older version software LBC restart

06:35 Fast OB VV to test Tech Chips (RH = 92.7% !)

No Zernike polymonials appeared on the LBC log ! Same problem

Seeing is 1.4 arcsec V-band

Vincenzo switched back to the latest software version and calling A.Di Paola

06:43 Slewing to El=+84.5 and DOFPIA to check collimation Table

LBCB conmverged in 2 iterations! LBCR iterations:

-590 -13 -279 149 -187 -202 -21 0.74 0.31
-1464 540 630 -1050 -912 553 10 0.71 0.95
4775 2422 -3586 2831 1824 -447 100 0.57 2.74
-4728 -2464 3804 -1351 -1242 406 -2 0.70 2.45

Red oscillates and diverges. Stop DOFPIA and Clear Active Optics on RED

06:56 Again DOFPIA

Blue converged in 3 iterations Red converged in 7 iterations

lbcb.20090914.070803.fits lbcr.20090914.070801.fits

07:11 Slewing down to EL=+35 and DOFPIA to check Collimation Table

Blue lbcb.20090914.072354.fits
-178 -3 -366 42 0 24 -4 1.17 0.15

Red lbcr.20090914.072356.fits
1 -224 -311 -138 -108 155 -16 0.87 0.19

07:26 Take 60s images VV to check IQ at EL=+34

lbcb.20090914.072622.fits FWHM =1.1"
lbcr.20090914.072620.fits FWHM=1.3"

Seeing looks good, slew to EL=+85 to see if its better and I can measure the Filters Focus Offset.

07:35 Now EL=+85 and DOFPIA (T_air = 8.3 RH=88.5% T_mirror=10.5)

Blue lbcb.20090914.074428.fits
-81 117 -111 -33 -163 -9 8 1.09 0.12

Red lbcr.20090914.074426.fits
56 113 12 -78 -101 -42 -21 0.44 0.11

07:47 check IQ with two 60s images VV

lbcb.20090914.074825.fits fwhm = 0.98"
lbcr.20090914.074823.fits fwhm = 1.98" (some issues? stars little elongated)

Repeat the test

lbcb.20090914.075405.fits fwhm = 1.12"
lbcr.20090914.075402.fits fwhm = 1.94"

Red is the same.

Recollimate

Note: A. di Paola is not yet available to work on the software but will be shortly.

Blue lbcb.20090914.075934.fits 198 -172 323 -233 -148 17 4 0.84 0.20
Red lbcr.20090914.075932.fits 17 69 111 -332 216 13 9 0.42 0.20

08:01 check IQ with two 60s images VV

lbcb.20090914.080205.fits fwhm=0.97"
lbcr.20090914.080205.fits fwhm=1.45"

Red still worse than Blue even if DOFPIA converged well.

08:07 Test Guide at EL=+81

lbcb.20090914.080829.fits fwhm=1.0"
lbcr.20090914.080205.fits fwhm=1.34"

Guide looks ok but red is still more elongated.

08:28 Red is too freaky to measure focus offsets so I give the Telescope to the LBC team

Now: RH=93.4% T_air=7.8

08:35 Close Dome RH jumped to 97.8% (My opinion: we should close at 90% maximum; 95% sounds like an amateur observatory limit -:)

09:50 Test Binning 2x2 (dome still closed)

lbcb.20090914.095001.fits lbcr.20090914.095001.fits

Ok it works

_Note by OPK 16 September 2009: I looked at the 2x2 binned images. LBCB image: lbcb.20090914.095001.fits is completely saturated. This was the problem we used to have when trying to bin. The LBC RED image, however, is OK - not saturated. So it looks like binning works for Red but not for Blue. But, now that the Red controller was swapped out for the spare, does binning with Red still work?_

10:00 Re-open

Vincenzo does some Sci/Commissioning INAF OBs

-- MarcoPedani - 14 Sep 2009
I Attachment Action Size Date Who Comment
daq.20090914.gifgif daq.20090914.gif manage 47 K 15 Sep 2009 - 00:28 JohnHill Plot of Telescope and Mirror Temperatures
Topic revision: r4 - 16 Sep 2009, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback